$SPAD/src/input patch51.input

Timothy Daly December 3, 2016

Abstract

These are bug fixes available in patch51

Contents

2

— * —

)set break resume
)spool patch51.output
)set message test on
)set message auto off
)clear all

———-

The besselK function was missing a minus sign as of patch50. This is fixed as of patch51. — * —

--S 1 of 1 bug #355 fix
D(besselK(a,x),x)
--R
--R
--R        - besselK(a + 1,x) - besselK(a - 1,x)
--R   (1)  -------------------------------------
--R                          2
--R                                                    Type: Expression(Integer)
--E 1
)spool
)lisp (bye)

———-

References

[1] Milton Abramowitz and Irene A. Stegun, eds. (1965). Handbook of Mathematical Functions with Formulas, Graphs, and Mathematical Tables. New York: Dover. ISBN 0-486-61272-4, Equations 9.1.27 and 9.6.26.